Find the sum of the first 220 positive even integers.
A)48,180
B)48,620
C)48,841
D)48,400

Answer by mananth(16946)

a=2,d=2
n=220
Sn=n/2[2a+(n-1)d]

=220/2[2*2+(220-1)*2]

=110
=110*442
=48620